From nch@roe.ac.uk Fri Apr 25 15:10:19 2008 Date: Fri, 25 Apr 2008 14:39:33 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Mark Holliman , Nigel Hambly , Nicholas Cross , Rob Blake , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Brian Walshe , Jim Emerson , Malcolm Stewart , Lorenzo Rimoldini , Mike Irwin , Peredur Williams , Bob Mann , Stephen Warren Subject: WFAU VDFS Science Archive weekly project meeting minutes, 25/04/08 Minutes of WFAU VDFS Science Archive meeting: 25th April 2008 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, ETWS, PMW, RSC, RPB, MAR, MSH, LGR, BCW Apologies: JPE, JMS, AL, RGM, NJC DONM: 10am, Friday 2nd May 2008 in the Vista Hut Actions discharged this week: ----------------------------- ACTION: RPB to finish restoring Ahmose and get it back to a usable condition, hopefully by early next week. Discharged ACTION: RPB to ensure all unnecessary logins are disabled on archive servers. Discharged; done on both Windows and Linux systems. Actions partly discharged but continuing: ----------------------------------------- ACTION: MSH to liaise with IT Support over set up of NFS file share from hatshepsut Continues; awaiting response from ITSG (ticket needed to be resubmitted) Actions carried forward from 18/04/08 meeting: ---------------------------------------------- ACTION: RPB to convene a meeting of interested parties to discuss locations and importance of all data help on WFAU servers to organise a consistent backup and DR strategy. Continues Specific points and new actions: -------------------------------- Project management: Nothing this week. WFCAM & VISTA updates: No news this week. Comments and issues arising from CASU fortnightly minutes: No new minutes this week. Networking: Following on from the restoration of the ingest DB on load server ahmose, there was some discussion as to the status of XMM/2XMM. ACTION: RSC to check out XMM/2XMM. Regarding the copying of BestDR5 to hatshepsut for the all-sky SSA, RPB noted that this is still proving difficult because of it's size, limitations in Symantec BackupExec, and the wish to minimise downtime of this much-used database. Work will pause while the present host server is upgraded to SQL2005 to see if this helps out in any way. WSA Operations: NCH thanked RPB for restoring the ingest DB so promptly following the recent hardware failures. In advance of the major recalibration the team agreed that a monthly full backup should be done this weekend. ACTION: RPB to make a full ingest DB (ahmose.WSA) backup this weekend ETWS noted: - Updated WSA with new filenames for the data moved off khafre and given IT support the all clear to investigate khafre. Hardware NCH raised the issue of the distribution of large "external" databases on the public servers, and what our future strategy would be in the light of storage requirements, cross-querying etc. MAR also pointed out potential trawl speed issues for networked storage resources. Solutions involving NAS were discussed at length, but in the short term the decision was made to stick with the original plan, and upgrade public server thutmose to SQL2005 to then investigate any new features like mirroring, copying and distributed query performance between the two SQL2005 instances hosted on thutmose and hatshepsut. ACTION: RPB to backup master and other system DBs on thutmose prior to upgrade to SQL2005. ACTION: RPB to upgrade thutmose to SQL2005 on Monday; MSH and MAR to note downtime in the usual ways in the user interfaces. Software: ETWS reported: - Finished implementation of illumination correction. - Wrote scripts for graphical testing of implemented corrections and started testing phase. Regarding testing of the illumination correction, NJC identified some useful fields for test purposes; NJC also noted that the switchable corrections needed to be driven by file type, and ETWS has implemented this. NCH thanked ETWS for his hard work on implementing the illumination correction, and also noted that once we have convinced ourselves that it is working correctly, Annette Ferguson and student Aram Kang have offered to check the photometry in their programme U/05B/7 against a catalogue product supplied and calibrated by MJI to make sure that we all agree (e.g. that we have all the signs the right way around!). This is a top priority for the next week. RSC noted that NJC's new prescription for boundary quality bit flagging could do with a look-over. ACTION: RSC and NCH to check NJC's new boundary quality error bit flag algorithm. Otherwise, RSC noted that he has completed software changes for schema and is presently testing it all thoroughly. Survey Data Release: MAR noted that QC1 for DR4 has progress very well, and should be closed out sometime next week. Non-survey Data Release: RSC noted that script production of WFCAMOPENTIME (and also WFCAMPROPRIETARY as a fail-over for the loadserver ingest DB) is now fully automated. ACTION: RSC to create open time/proprietary metadata DBs on public server Astrogrid deployment: MSH noted that the new Astrogrid components are now up and running bar a few teething troubles; MAR asked that some usernames for the group be added to the new community. ACTION: MSH to set up usernames for MAR, LGR and NCH on the new AG community Miscellaneous: MSH noted that he will be attending the VO Workshop in Munich in June http://www.euro-vo.org/dcaworkshop2008/ ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. =============================================================